OpenStudy (anonymous):

Please check these two algebra questions! :) simplify x^2+8x/x^2 A) x+8/x^2 B) x+8/x C) 8x D) 8 I think it's D simplify. x^2-16x+64/x2-64 A) -1 B) x-1/x+1 C) x-8/x+8 D) x-8/x^2-64 I started working it out and got stuck with -16x+64/-64. I'm not sure what to do next.

OpenStudy (raden):

1) u need factor out to simpify it (x^2+8x)/x^2 = x(x+8)/x^2 = ... 2) (x^2-16x+64)/(x^2-64) = (x-8)(x-8)/(x-8)(x+8) = ... (cancel out the similar terms)